### v1

# Does a Rising Tide Lift All Boats? Bias Mitigation for AI-based CMR Segmentation

Abstract
Artificial intelligence (AI) is increasingly being used for medical imaging tasks. However, there can be biases in the resulting models, particularly when they were trained using imbalanced training datasets. One such example has been the strong race bias effect in cardiac magnetic resonance (CMR) image segmentation models. Although this phenomenon has been reported in a number of publications, little is known about the effectiveness of bias mitigation algorithms in this domain. We aim to investigate the impact of common bias mitigation methods to address bias between Black and White subjects in AI-based CMR segmentation models. Specifically, we use oversampling, importance reweighing and Group DRO as well as combinations of these techniques to mitigate the race bias. Furthermore, motivated by recent findings on the root causes of AI-based CMR segmentation bias, we evaluate the same methods using models trained and evaluated on cropped CMR images. We find that bias can be mitigated using oversampling, significantly improving performance for the underrepresented Black subjects whilst not significantly reducing the majority White subjects’ performance. Group DRO also improves performance for Black subjects but not significantly, while reweighing decreases performance for Black subjects. Using a combination of oversampling and Group DRO also improves performance for Black subjects but not significantly. Using cropped images increases performance for both races and reduces the bias, whilst adding oversampling as a bias mitigation technique with cropped images reduces the bias further.
Keywords:
Bias mitigation CMR imaging Segmentation.
1 Introduction
Artificial intelligence (AI) is increasingly being used to aid medical diagnosis, prognosis and treatment planning. However, AI models have been shown to exhibit bias by protected attributes in many different applications. For example, in cardiac magnetic resonance (CMR) image segmentation, both sex [1] and race [2, 3, 4] bias have been reported. Sex bias has also been found in chest X-ray image classification [5, 6]. AI bias can have detrimental downstream impacts in medical imaging applications. For example, CMR segmentations are used to derive biomarkers whose values impact patient management, so greater errors in these biomarkers for certain protected groups can lead to inappropriate treatment choices and worse outcomes [4].
Previous work has aimed to address biases in AI models for medical imaging tasks by using bias mitigation methods. [7] evaluated and mitigated bias in chest X-ray classification models using methods based on oversampling minority protected groups and Pareto minimax optimisation. [8] proposed a framework with eleven algorithms which aimed to measure and mitigate biases in medical imaging classification datasets. [9] evaluated stratified batch sampling, a balanced dataset model and a protected group-specific model for orthopaedic image segmentation. In CMR segmentation, [3] proposed oversampling of minority protected groups and a multi-task learning approach which learnt both a segmentation model and a protected attribute classifier.
Recently, an investigation was performed to discover the root cause(s) of AI-based CMR segmentation bias [10]. One key finding of this work was that distributional differences in areas outside the heart were significant in the learning of biased representations in AI segmentation models. This finding opens up new avenues of enquiry in bias mitigation, which we explore in this paper.
The contribution of this paper is to perform the most extensive and comprehensive investigation to date of bias mitigation methods in AI-based CMR segmentation. We investigate multiple mitigation methods as well as combinations of these methods. We also investigate AI models trained using CMR images that are cropped to remove features outside of the heart that may be leading to bias, and evaluate the effectiveness of bias mitigation techniques in this setting. Our results show that CMR segmentation bias can be mitigated to produce results that have high accuracy but are not significantly different between races.
2 Methods
We use CMR images from the UK Biobank [11]. The dataset consists of end diastolic (ED) and end systolic (ES) cine short-axis images from 5,778 subjects. Manual segmentation of the left ventricular blood pool (LVBP), left ventricular myocardium (LVM), and right ventricular blood pool (RVBP) was performed for the ED and ES images of each subject. The LV endocardial and epicardial borders and the RV endocardial border were outlined using cvi42 (version 5.1.1, Circle Cardiovascular Imaging Inc., Calgary, Alberta, Canada). The same guidelines were provided to a panel of ten experts with one expert annotating each image. Each expert was provided with a random selection of images for annotation which included subjects of different sexes and races. They were not provided with demographic information about the subjects.
Previous work [2, 1] has shown that bias is greater when the imbalance between races in the training set is greater. Therefore, we curate a dataset where biases would be significant to allow for better evaluation of mitigation methods. Our training set comprises 15 Black subjects and 4,221 White subjects.
2.1 Baseline Model
We trained a baseline 2D nnU-Net v1 model [12] to segment the LVM, LVBP and RVBP using the training parameters from [1]. The demographic information of the subjects used in the training and testing datasets can be seen in Table 1 and Table 2. Training was performed using only Black and White subjects but testing was performed using all other available subjects including Mixed, Asian, Chinese and Other.
Table 1: Characteristics of subjects used in the training dataset. Mean (standard deviation) values are presented for each characteristic. Statistically significant differences between subject groups and the overall average are indicated with an asterisk * (p $<$ 0.05) and were determined using a two-tailed Student’s t-test.
Health measure Overall White Black # subjects 4236 4221 15 Age (years 64.6 (7.7) 64.6 (7.7) 57.0 (4.9)* Weight (kg) 77.0 (15.0) 77.0 (15.0) 74.8 (11.9)* Standing height (cm) 169.5 (9.2) 169.5 (9.2) 168.8 (6.9)* Body Mass Index 26.7 (4.3) 26.7 (4.3) 26.2 (3.7)*
Table 2: Characteristics of subjects used in the test dataset. Mean (standard deviation) values are presented for each characteristic. Statistically significant differences between subject groups and the overall average are indicated with an asterisk * (p $<$ 0.05) and were determined using a two-tailed Student’s t-test.
Health measure Overall White Mixed Asian Black Chinese Other # subjects 1542 469 170 387 223 111 182 Age 61.7 (7.9) 64.8 (7.7)* 59.8 (7.2)* 61.0 (8.2)* 59.1 (7.1)* 59.7 (6.4)* 62.1 (7.5) Standing height 167.4 (9.2) 169.8 (9.4)* 166.4 (8.5) 166.7 (8.6) 168.7 (9.4) 162.4 (7.1)* 165.4 (9.7)* Weight 75.1 (15.1) 77.7 (15.1)* 75.1 (15.3) 72.3 (12.8)* 82.0 (16.0)* 63.5 (9.9)* 73.3 (15.5) Body Mass Index 26.7 (4.4) 26.9 (4.2) 27.1 (5.2) 25.9 (3.7)* 28.8 (5.1)* 24.0 (3.1)* 26.6 (4.4)
2.2 Oversampling
We also trained a 2D nnU-Net model using the same data as the baseline but applied oversampling during batch selection [13]. Oversampling refers to the process of increasing the sampling of a minority group in a dataset. Here, we oversample the Black subjects in the training dataset so that they were equal to the number of White subjects in each batch used during training. This was performed using random sampling with replacement so each subject could in principle be selected more than once in a training batch. The same training parameters from [1] were used.
2.3 Reweighing
An nnU-Net model was also trained using a reweighing mitigation strategy [13]. Reweighing refers to the process of increasing the importance of under-represented groups to the model. This can be done as a pre-processing step or during training by adapting the loss function. We performed this during training by adding a weighting term to the combined Cross Entropy (CE)-Dice loss function of the nnU-Net. Each group was weighted inversely proportionally to the group size, as shown in Eq. 2. The weights were then normalised so that they summed to 1. The training parameters from [1] were again used.
Weights per group: $\quad w_{g}=\frac{n_{G}}{n_{g}+\epsilon},\quad\text{where }\epsilon=10^{-6}.$ (1)
Normalized weights: $\quad\hat{w}_{g}=\frac{w_{g}}{\sum_{j=1}^{N_{G}}w_{j}},\quad g=1,2,\dots,N_{G}.$ (2)
where $n_{g}$ is the number of samples in protected group $g$, $n_{G}$ is the number of samples in all groups and $N_{G}$ is the number of groups.
2.4 Group Distributionally Robust Optimisation
The final mitigation approach was Group Distributionally Robust Optimisation, or Group DRO, which was first proposed in [14]. The method aims to optimise the performance of the worst-performing group in a dataset. The Group DRO loss function can be formalised as:
${L_{DRO}}=max_{g\in G}\frac{1}{{{n_{g}}}}\sum\nolimits_{i\in g}{L\left(y_{i},\hat{y_{i}}\right)}$ (2)
where $L$ is the loss function computed between predicted labels $\hat{y_{i}}$ and ground truth labels $y_{i}$.
In this method, CE loss was used instead of CE-Dice loss which was used for the oversampling and reweighing experiments. Group DRO uses losses from individual samples to calculate the average loss for the groups. However, Dice loss is calculated using global statistics of the true positives, false positives and false negatives for a group or batch. It is non-additive as the numerator and denominator will change if the calculation is performed on a per-sample basis rather than for a group or batch. For Dice loss, the average group loss and global loss are different, which causes instability in training.
The code for the bias mitigation models is available at https://github.com/tiarnaleeKCL/nnUNet-bias-mitigation
2.5 Training Using Cropped Images
Following the findings of [10], which found that areas outside the heart were a contributing factor to CMR segmentation performance bias, we also performed experiments using all of the above techniques for a nnU-Net model trained using cropped CMR images. The images were cropped around the heart using a bounding box defined based on the ground truth segmentation. All images were cropped to the same size, i.e. the size of the largest heart in the dataset.
2.6 Training Using Combinations of Mitigation Methods
The final experiment combined the mitigation methods into pairs to test whether combinations of methods would improve performance. This results in three additional methods: oversampling + Group DRO, reweighing + Group DRO, and oversampling + reweighing. These methods were applied in the same way as above. CE loss was used for experiments with Group DRO and CE-Dice loss was used for oversampling + reweighing.
2.7 Metrics
For each of the methods described above, performance was measured by finding the overall Dice similarity coefficient (DSC) for subjects in the test set. Performance was measured using the median and inter-quartile range (IQR) of the DSC. Fairness metrics are also reported: the fairness gap (FG) as defined in [1] and skewed error ratio (SER) as defined in [3]. The SER is defined as $SER=\frac{max_{g}(1-D_{g})}{min_{g}(1-D_{g})}$ where $D_{g}$ is the median DSC for protected group $g$. This measures the ratio of the errors between the median DSCs for the worst-performaing and best-performing protected groups. A more fair model will have a FG closer to 0 and a SER closer to 1.
3 Results
The results of the baseline and the three bias mitigation methods can be seen in Table 3 and in Fig. 1. Oversampling was the only method to increase performance for the Black subjects such that there was no significant difference between the median DSC scores of the Black and White subjects. The method increased the median performance for Black subjects by 0.045. Fairness performance metrics (SER and FG) also decreased, showing more equitable performance. Using oversampling also caused performance to increase for the other races compared to the baseline, as can be seen in Fig. 1(a) and Fig. 1(b). This is perhaps surprising as performance on these other races was not optimised during training.
The other mitigation methods did not significantly improve performance for the Black subjects. Reweighing resulted in worse performance for the Black subjects, with the median DSC decreasing. Group DRO resulted in increased performance for the Black subjects but performance remained significantly lower than for the White subjects. Both oversampling and Group DRO slightly decreased median DSC for White subjects (although not statistically significantly), but reweighing significantly decreased median DSC for White subjects.
Table 3: DSC values for each of the bias mitigation methods. The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test. * p$<$ 0.0001. The best median DSC score for Black and White subjects is shown in bold.
Baseline * Oversampling p = 0.22 Reweighing * Group DRO * White Black White Black White Black White Black Median 0.896 0.846 0.894 0.891 0.891 0.831 0.893 0.865 IQR 0.046 0.064 0.045 0.049 0.049 0.069 0.047 0.051 SER 1.486 1.032 1.544 1.260 Fairness gap 0.050 0.003 0.059 0.028
Baseline
Oversampling
Reweighing
Group DRO
Figure 1: Overall DSC for bias mitigation methods on uncropped images. The dashed line indicates median DSC for White test subjects.
As shown in Table 4, the baseline model trained using cropped images improved performance for both White and Black subjects and reduced bias. Oversampling significantly improved performance for the Black subjects compared to the baseline, resulting in performance that was higher than for White subjects. Group DRO improved the DSC for Black subjects but this increase was not significant. Reweighing made performance slightly worse for both races, with median DSC decreasing and SER and FG measures increasing.
Table 4: DSC values for each of the bias mitigation methods using cropped images. The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test. * p$<$ 0.0001. The best median DSC score for Black and White subjects is shown in bold.
Baseline * Oversampling * Reweighing * Group DRO * White Black White Black White Black White Black Median 0.919 0.893 0.913 0.922 0.916 0.887 0.918 0.908 IQR 0.032 0.041 0.030 0.039 0.033 0.043 0.033 0.038 SER 1.326 1.115 1.352 1.130 Fairness gap 0.026 --0.009 0.029 0.011
Finally, combining the mitigation methods did not produce significantly less biased results, as shown in Table 5. All three combinations decreased performance for White subjects but increased performance for Black subjects. The best combination was oversampling and Group DRO which reduced the performance for the White subjects the least and improved performance for the Black subjects the most, leading to the lowest FG.
Table 5: DSC values for the combined bias mitigation methods using original sized images. The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test. * p$<$ 0.0001. The best median DSC score for Black and White subjects is shown in bold.
Baseline Oversampling + Group DRO Group DRO + Reweighing Oversampling + Reweighing White Black White Black White Black White Black Median 0.896 0.846 0.889 0.881 0.893 0.863 0.889 0.871 IQR 0.046 0.064 0.048 0.054 0.047 0.055 0.051 0.057 SER 1.486 1.08 1.285 1.161 Fairness gap 0.050 0.009 0.031 0.017
4 Discussion
This work has performed a comprehensive examination of bias mitigation methods for AI segmentation models used for cine CMR images. We have shown that bias in CMR segmentation models can be mitigated by using such methods. In particular, oversampling minority subjects reduces bias so that there is no significant difference between the performance of the Black and White subjects. Although oversampling did not add any extra information to the dataset, the method allowed the network to train on Black subjects more frequently than if oversampling was not used, allowing for better balance between protected groups. It could be anticipated that training using a small number of (oversampled) Black subjects would increase the risk of overfitting to those subjects, having a detrimental effect on generalisation, but this effect was not seen in our experiments as test performance remained high. Previous work in [10] showed that race could be classified from cine CMR images, indicating that there are distinct features in the images of different races that are recognisable to AI models. Using oversampling will allow the network to see more of these distinct features to learn better representations of the under-represented group.
Reweighing did not improve segmentation performance, instead decreasing performance for both protected groups. This may be due to increased importance being given to a small group of subjects, decreasing focus on the larger group. As described in [2] and [1], when White subjects comprised 75% of the training set, their segmentation performance was still lower than for the Black subjects who comprised 25% of the training set. This suggests that segmentation of the White subjects’ images may be a more difficult task as there may be more outliers and variation in the hearts than in the Black subjects. Reweighing these White subjects so that their importance is lower in the loss function could decrease accuracy.
Using cropped images increased performance and reduced bias compared to using uncropped images. Using oversampling with cropped images reduced bias further, but interestingly not to the same extent as using uncropped images. Also, note that to use a model trained using cropped images in clinical practice, a region-of-interest proposal network would be required to identify the cropping region at test time [15].
The combination of Group DRO and oversampling produced a model which improved performance for Black subjects but was still significantly lower than for White subjects. Therefore, using a combined method was better than using Group DRO alone but worse than oversampling alone.
This work has some limitations. For example, only Black and White subjects were used during training to control for other factors which may affect mitigation methods. Future work could investigate the effect of mitigating biases using multiple races or in other protected attributes such as age and socioeconomic status. It would also be useful to investigate the performance of bias mitigation strategies under external validation.
5 Acknowledgements
This work was supported by the Engineering & Physical Sciences Research Council Doctoral Training Partnership (EPSRC DTP) grant EP/T517963/1. This research has been conducted using the UK Biobank Resource under Application Number 17806